An airplane travels 1800 km in 6 h flying with the wind. It travels only two thirds as far in the same time against the wind. Find the speed of the wind.
Thank you~

Answer by mananth(16946)

plane speed = 300 mph
wind speed = x mph
wind speed
speed against wind 300 - x
speed with wind 300 + x

Distance against wind 1200 miles
Distance with wind 1800 miles

t=d/r time against = time with wind
1200/(300-x)=1800/(300+x)
1200(300+x)=1800(300-x)
360000 + 1200x = 540000-1800x
1800x +1200x = 180000
3000 x = 180000
/ 3000
x = 60 mph wind speed
m.ananth@hotmail.ca